# Amortized Bayesian Meta-Learning for Low-Rank Adaptation of Large Language Models

### v2

@@ Abstract
Fine-tuning large language models (LLMs) with low-rank adaptation (LoRA) is a cost-effective way to incorporate information from a specific dataset. However, it is often unclear how well the fine-tuned LLM will generalize, i.e., how well it will perform on unseen datasets. Methods have been proposed to improve generalization by optimizing in-context prompts, or by using meta-learning to fine-tune LLMs. However, these methods are expensive in memory and computation, requiring either long-context prompts or saving copies of parameters and using second-order gradient updates. To address these challenges, we propose Amortized Bayesian Meta-Learning for LoRA (ABMLL). This method builds on amortized Bayesian meta-learning for smaller models, adapting this approach to LLMs while maintaining its computational efficiency. We reframe task-specific and global parameters in the context of LoRA and use a new hyperparameter to balance reconstruction accuracy and the fidelity of task-specific parameters to the global ones. ABMLL provides effective generalization and scales to large models such as Llama3-8B. Furthermore, as a result of using a Bayesian framework, ABMLL provides improved uncertainty quantification. We test ABMLL on CrossFit and Unified-QA datasets and find that it outperforms existing methods on these benchmarks in terms of both accuracy and expected calibration error.
@@ 1 Introduction
Large language models (LLMs) handle a variety of tasks reasonably well (radford2019language). However, to tailor LLMs to specific domains, fine-tuning on specific datasets is often necessary. While methods such as low-rank adaptation (LoRA; Hu2021LoRALA) fine-tune a pretrained LLM cost-effectively, a fine-tuned LLM is limited to the domain it is trained on. Its performance may not improve in other domains and sometimes worsens as it suffers from catastrophic forgetting. Such catastrophic forgetting may result in overfitting and erasing existing capabilities of the pretrained LLM (Lazaridou2021MindTG; luo2025empiricalstudycatastrophicforgetting).
@@
Meta-learning is a strategy for solving this problem, training models on a variety of tasks in a way that supports generalization across tasks (Finn2017ModelAgnosticMF). However, meta-learning typically requires a large amount of computation and memory, making it challenging to apply to LLMs. One form of meta-learning that has been applied to LLMs involves fine-tuning models on in-context prompt-response examples (min-etal-2022-metaicl; chen-etal-2022-meta). Another more traditional approach, MAML-en-LLM (mamlenllm), adapts the Model-Agnostic Meta-Learning (MAML) (Finn2017ModelAgnosticMF) framework to LLMs. However, both methods are limited in the size of the language models that can be used: the former requires long-context prompts, whereas the latter uses second-order gradient updates and saves a model for each task.
@@
Recent work on Amortized Bayesian Meta-Learning (ABML; ravi2018amortized) addresses some of the computation and memory requirements of meta-learning. This approach posits a generative model over parameters where task-specific parameters are generated from global parameters, and inference over task-specific parameters is amortized. In other words, the conditional distribution over task-specific parameters is shared across tasks, implying that computation and memory costs stay constant with respect to the number of tasks. This approach thus offers a path towards efficient meta-learning for LLMs. However, several challenges exist. First, we need to specify the generative model over the weight space for LLMs. Second, the enormous size of LLMs makes training difficult, as the scale of probabilities assigned to the model variables can overwhelm the influence of the data likelihood.
@@
In this paper, we present a solution to these problems, making it possible to take a Bayesian approach to fine-tuning LLMs using ABML (see Figure 1). To define the underlying generative model and efficiently characterize the distributions involved, we use LoRA to express both the model weights and their uncertainty. We introduce a new prior over global variables that accounts for the spread of the parameters learned in the pretrained model. We also introduce an adjustable hyperparameter that balances reconstruction accuracy and the fidelity of task-specific parameters to global values.
Using amortized Bayesian meta-learning for LLM fine-tuning, we achieve significantly stronger performance on unseen tasks compared with regular LoRA fine-tuning. We show that amortized Bayesian meta-learning results in fine-tuned LLMs that are accurate on domain-specific tasks, more generalizable to new tasks, and provide better uncertainty estimation. Our method avoids the computation and memory overhead of other meta-learning approaches, making it adaptable to larger models such as Llama3-8B (grattafiori2024llama3herdmodels) with minimal memory increase from regular LoRA. Finally, because one advantage of Bayesian methods is a natural regularization, we prune fine-tuned LLMs and show that our method is significantly stronger under pruning than both regular fine-tuning and other meta-learning methods.
@@ 2 Related Work
Meta-learning methods in LLMs.
Extensive work has explored meta-learning as a method for improving generalization in machine learning systems, although these approaches were typically developed for models in the pre-LLM era (Finn2017ModelAgnosticMF; snell-prototypical; ravi2018amortized; Nichol2018OnFM). mamlenllm adapted Model-Agnostic Meta-Learning (Finn2017ModelAgnosticMF) to LLMs. However, this adaptation is more expensive in computation and memory than our method, requiring second-order gradient updates and saving a model for each task. More recently, lift2025 proposed a hierarchical Bayesian approach to LoRA meta-learning, but its parameters also increase linearly with number of tasks. As a result, we are able to evaluate performance on larger models than the ones tried in these two papers.
In a different approach, min-etal-2022-metaicl and chen-etal-2022-meta explored meta-learning for LLMs using in-context learning. These works show that it is possible to fine-tune LLMs on in-context examples and achieve generalization. However, our approach does not require curation of such examples, does not place constraints on the size of the context window of a model, and is more scalable.
@@
Approaches to capturing uncertainty for LLMs can rely on the intrinsic representation of uncertainty in the model or focus on capturing extrinsic uncertainty about model parameters. Intrinsic approaches produce better uncertainty calibration via prompt engineering and sampling (gruver-time-series) or learning an external model (pmlr-v235-shen24c). Extrinsic approaches include using fine-tuning methods to incorporate uncertainty, such as training LoRA with ensembles (Balabanov2024UncertaintyQI), Laplace approximation (Yang2023BayesianLA), and variational inference (wang2024blob). Our work takes the extrinsic approach but differs from existing approaches by using meta-learning to achieve generalization across datasets.
@@ 3.1 Low-Rank Adaptation (LoRA)
LoRA (Hu2021LoRALA) fine-tunes LLM weights on a low-rank space to improve efficiency compared with regular fine-tuning. Let $\mathbf{W}_{0}$ of size $d_{\text{out}}$-by-$d_{\text{in}}$ denote a weight matrix from a pretrained LLM. Let $\mathbf{x}$ denote the input to $\mathbf{W}_{0}$, and $\mathbf{z}$ denote the output of $\mathbf{W}_{0}$, LoRA fine-tunes the pretrained weight $\mathbf{W}_{0}$ by adding a low-rank matrix comprised of two trainable matrices,
@@ 3.2 Approaches to Meta-Learning
Meta-learning aims to find a set of initial model parameters that can be rapidly adapted to new, unseen tasks with a few gradient steps (schmidhuber1987evolutionary; bengio1991learning; caruana1998multitask). The strategy for doing so is to generalize from the shared statistical structure across tasks: by extracting this structure, a model can “learn to learn.” A common setting in which meta-learning is used is few-shot learning, where each task might only provide a small number of examples but there are many such tasks. Personalizing large language models through modification of their weights, rather than their prompts or the context they condition on, is a natural setting for using this approach, where each user might only have a limited amount of data available but a group of users may have similar interests.
@@
A popular approach to meta-learning is the Model-Agnostic Meta-Learning (MAML; Finn2017ModelAgnosticMF) algorithm. This algorithm runs two training loops: an inner loop for task-specific adaptation and an outer loop for meta-optimization. Let $D_{i}$ denote a batch of data from task $i$, $p_{\theta}$ denote the prediction model parameterized by $\theta$, and $\alpha,\beta$ denote gradient descent step sizes. The goal of learning is to obtain a set of parameters $\theta_{i}$ for each task and a global set of parameters $\theta$ that are used to initialize learning for all tasks. In a given epoch, for each task $i$, MAML conducts an inner loop gradient update,
@@ 3.3 Amortized Bayesian Meta-Learning
Amortized Bayesian Meta-Learning (ABML; ravi2018amortized) improves upon MAML-based meta-learning frameworks by representing uncertainty with a Bayesian approach. It also amortizes inference over the parameters so that memory no longer increases linearly with the number of tasks.
@@ 4 Method
ABMLL uses the the objective of Eq. 1 from ABML. In our setting, $\theta$ and $\phi_{i}$ are the global and task-specific model parameters produced as the output of LoRA adapters. On a high level, the generative process is
$\theta$ $\sim p(\theta),$
$\phi_{i}$ $\sim p(\phi_{i}|\theta),$
$D_{i}$ $\sim\text{LLM}(\phi_{i}),$
@@
where $i$ represents any task $i$, and $\text{LLM}(\phi_{i})$ denotes the LLM considered as a probabilistic model that takes $\phi_{i}$ as its inputs and outputs token sequences with joint probabilities defined by the LLM’s autoregressive predictive distribution. By positing that task specific variables $\phi_{i}$ are generated from global variables $\theta$, the model is encouraged to learn a generalizable space of parameters with fast adaptions to different tasks. We provide pseudocode in Algorithm 1 to illustrate our approach. For any LLM layer with pretrained weights $\mathbf{W_{0}}$, the quantities for our extension to ABML are:
$\mu_{\theta}$ $=\mathbf{B}_{\mu_{\theta}}\mathbf{A}_{\mu_{\theta}},$
$\log\sigma_{\theta}^{2}$ $=\mathbf{B}_{\sigma_{\theta}}\mathbf{A}_{\sigma_{\theta}}+c\mathbf{I},$
$\mu_{\phi}$ $=\mathbf{B}_{\mu_{\phi}}\mathbf{A}_{\mu_{\phi}},$
$\log\sigma_{\phi}^{2}$ $=\mathbf{B}_{\sigma_{\phi}}\mathbf{A}_{\sigma_{\phi}}+c\mathbf{I},$
$p(\phi_{i}|\theta)$ $=\mathcal{N}(\phi_{i};\mu_{\theta}+\mathbf{W_{0}},\sigma_{\theta}^{2}),$
$\hskip 0.0ptq_{\theta}(\phi_{i}|D_{i})$ $=\mathcal{N}(\phi_{i};\mu_{\phi}+\mathbf{W_{0}},\sigma_{\phi}^{2}),$

### v3

# Meta-Learning at Scale for Large Language Models via Low-Rank Amortized Bayesian Meta-Learning

@@ Abstract
Fine-tuning large language models (LLMs) with low-rank adaptation (LoRA) is a cost-effective way to incorporate information from a specific dataset. However, when a problem requires incorporating information from multiple datasets – as in few shot learning – generalization across datasets can be limited, driving up training costs. As a consequence, other approaches such as in-context learning are typically used in this setting. To address this challenge, we introduce an efficient method for adapting the weights of LLMs to multiple distributions, Amortized Bayesian Meta-Learning for LoRA (ABMLL). This method builds on amortized Bayesian meta-learning for smaller models, adapting this approach to LLMs by reframing where local and global variables are defined in LoRA and using a new hyperparameter to balance reconstruction accuracy and the fidelity of task-specific parameters to the global ones. ABMLL supports effective generalization across datasets and scales to large models such as Llama3-8B and Qwen2-7B, outperforming existing methods on the CrossFit and Unified-QA datasets in terms of both accuracy and expected calibration error. We show that meta-learning can also be combined with in-context learning, resulting in further improvements in both these datasets and legal and chemistry applications.
@@ 1 Introduction
Large language models (LLMs) handle a variety of tasks well by default (Radford et al., 2019). However, fine-tuning is often necessary to tailor LLMs to specific domains. While methods such as low-rank adaptation (LoRA; Hu et al., 2021) fine-tune a pretrained LLM cost-effectively, the fine-tuned LLM is often limited to the domain it is trained on. Its performance may not improve in other domains and sometimes worsens because of catastrophic forgetting, which can erase existing capabilities (Lazaridou et al., 2021; Luo et al., 2023).
@@
In-context learning (ICL) adapts models to new tasks by conditioning on examples from the task in the prompt (Min et al., 2022; Chen et al., 2022). However, prompt-based approaches have limited expressivity, and their gains may diminish as the complexity of in-context information rises (Dou et al., 2026). In-weights meta-learning is another strategy for solving this problem, training models on a variety of tasks in a way that supports generalization across tasks (Finn et al., 2017). MAML-en-LLM (Sinha et al., 2024), adapts the Model-Agnostic Meta-Learning (MAML; Finn et al., 2017) framework to LLMs. However, this approach requires a large amount of computation and memory, using second-order gradient updates and saves a model for each task, making it challenging to apply at scale.
@@
We address this challenge by developing a meta-learning approach for LLMs that is practical to use at scale and can be combined with ICL. Reptile (Nichol et al., 2018) is a meta-learning algorithm that has constant scaling with respect to the number of tasks, but it does not retain the full expressivity of MAML. We propose another scalable meta-learning solution based on Amortized Bayesian Meta-Learning (ABML; Ravi and Beatson, 2019). This method posits a generative model over model weights where task-specific weights are generated from global weights. Conditional distributions over task-specific weights are shared across tasks, so computation and memory costs remain constant with respect to the number of tasks. This offers a path toward meta-learning for LLMs at scale, but introduces two challenges: specifying a generative model over LLM weight space, and making training stable when the scale of probabilities over model variables can overwhelm the data likelihood.
@@
We present a solution to these two problems (see Figure 1). To define the generative model and efficiently characterize uncertainty, we use LoRA to express both model weights and their uncertainty, with mean and variance parameters computed as LoRA outputs. We introduce a new prior over global variables that accounts for the spread of pretrained parameters, along with a hyperparameter that balances reconstruction accuracy and the fidelity of task-specific parameters to global values.
Using Amortized Bayesian Meta-Learning for LoRA (ABMLL), we show that meta-learning can work effectively at the scale of Llama3-8B and Qwen2-7B (Grattafiori et al., 2024; Yang et al., 2024) with only a small memory increase over regular LoRA. On the standard benchmarks CrossFit and UnifiedQA, ABMLL improves over regular LoRA in both accuracy and expected calibration error, while performing competitively with other scalable meta-learning baselines on unseen tasks. We further show that meta-learning can be combined with ICL, yielding additional gains on these datasets as well as in legal and chemistry applications. Finally, we show that ABMLL improves robustness under pruning relative to regular fine-tuning and other meta-learning methods, suggesting the promise of a Bayesian approach for scalable and robust adaptation.
@@ 2 Related Work
Generalization methods in LLMs.
Extensive work has explored meta-learning as a method for improving generalization in machine learning systems, although these approaches were typically developed for models in the pre-LLM era (Finn et al., 2017; Snell et al., 2017; Ravi and Beatson, 2019; Nichol et al., 2018). Sinha et al. (2024) adapted Model-Agnostic Meta-Learning (Finn et al., 2017) to LLMs. However, this adaptation is more expensive in computation and memory than our method, requiring second-order gradient updates and saving a model for each task. More recently, Kim and Hospedales (2025) proposed a hierarchical Bayesian approach to LoRA meta-learning, but its parameters also increase linearly with number of tasks. As a result, we are able to evaluate performance on larger models than the ones tried in these two papers.
In a different approach, Min et al. (2022) and Chen et al. (2022) explored meta-learning for LLMs using in-context learning (ICL). These works show that it is possible to fine-tune LLMs on in-context examples and achieve generalization. However, ICL has shown weaknesses in utilizing knowledge that are both novel and complex (Dou et al., 2026). We show that in-weights meta-learning can utilize ICL to achieve a better performance.
@@
Approaches to capturing uncertainty for LLMs can rely on the intrinsic representation of uncertainty in the model or focus on capturing extrinsic uncertainty about model parameters. Intrinsic approaches produce better uncertainty calibration via prompt engineering and sampling (Gruver et al., 2023) or learning an external model (Shen et al., 2024). Extrinsic approaches include using fine-tuning methods to incorporate uncertainty, such as training LoRA with ensembles (Balabanov and Linander, 2024), Laplace approximation (Yang et al., 2023), and variational inference (Wang et al., 2024). Our work takes the extrinsic approach but differs from existing approaches by using meta-learning to achieve generalization across datasets.
@@ 3.1 Low-Rank Adaptation (LoRA)
LoRA (Hu et al., 2021) fine-tunes LLM weights on a low-rank space to improve efficiency compared with regular fine-tuning. Let $\mathbf{W}_{0}$ of size $d_{\text{out}}$-by-$d_{\text{in}}$ denote a weight matrix from a pretrained LLM. Let $\mathbf{x}$ denote the input to $\mathbf{W}_{0}$, and $\mathbf{z}$ denote the output of $\mathbf{W}_{0}$, LoRA fine-tunes the pretrained weight $\mathbf{W}_{0}$ by adding a low-rank matrix that is the product of two trainable matrices,
@@ 3.2 Approaches to Meta-Learning
Meta-learning aims to find a set of initial model parameters that can be rapidly adapted to new, unseen tasks with a few gradient steps (Schmidhuber, 1987; Bengio et al., 1991; Caruana, 1998). The strategy for doing so is to generalize from the shared statistical structure across tasks: by extracting this structure, a model can “learn to learn.” A common setting for meta-learning is few-shot learning, where each task has few examples and there are many such tasks. Personalizing large language models through modification of their weights, rather than their prompts or the context they condition on, is a natural setting for using this approach, where each user might only have a limited amount of data available but a group of users may have similar interests.
@@
A popular approach to meta-learning is the Model-Agnostic Meta-Learning (MAML; Finn et al., 2017) algorithm. This algorithm runs two training loops: an inner loop for task-specific adaptation and an outer loop for meta-optimization. Let $D_{i}$ denote a batch of data from task $i$, $p_{\theta}$ denote the prediction model parameterized by $\theta$, and $\alpha,\beta$ denote gradient descent step sizes. The goal of learning is to obtain a set of parameters $\theta_{i}$ for each task and a global set of parameters $\theta$ that are used to initialize learning for all tasks. In a given epoch, for each task $i$, MAML conducts an inner loop gradient update,
@@ 3.3 Amortized Bayesian Meta-Learning
Amortized Bayesian Meta-Learning (ABML; Ravi and Beatson, 2019) improves upon MAML-based meta-learning frameworks by representing uncertainty with a Bayesian approach. It also amortizes inference over the parameters so that memory no longer increases linearly with the number of tasks.
@@ 4 Method
ABMLL uses the the objective of Eq. 1 from ABML. In our setting, $\theta$ and $\phi_{i}$ are the global and task-specific model parameters produced as the output of LoRA adapters. At a high level, the generative process is
$\theta$ $\sim p(\theta),\;\;\phi_{i}\sim p(\phi_{i}|\theta),\;\;D_{i}\sim\text{LLM}(\phi_{i}),$
@@
where $i$ indexes over tasks and $\text{LLM}(\phi_{i})$ denotes the LLM considered as a probabilistic model that takes $\phi_{i}$ as its inputs and outputs token sequences with joint probabilities defined by the LLM’s autoregressive predictive distribution. By positing that task-specific variables $\phi_{i}$ are generated from global variables $\theta$, the model is encouraged to learn a generalizable space of parameters with fast adaptions to different tasks. More detailed algorithm is given in the Appendix A.1, Algorithm 1. For any LLM layer with pretrained weights $\mathbf{W_{0}}$, the quantities for our extension to ABML are: